OpenStudy (anonymous):

a1=28, d= -3...... find a8

OpenStudy (diyadiya):

ok Formula ?

OpenStudy (anonymous):

a n =a 1 +(n−1)d a8=28+(8-1)-3

OpenStudy (diyadiya):

a8=28+(8-1)3

OpenStudy (anonymous):

its a -3 though.....

OpenStudy (diyadiya):

oh ok sorry a8=28+(8-1) x (-3)

OpenStudy (anonymous):

where does the 'x' come from ? ? ??

OpenStudy (diyadiya):

a n =a 1 +(n−1)d a n =a 1 +(n−1) x d We multiplied it remember ?

OpenStudy (anonymous):

oh i thought you meant the letter x

OpenStudy (diyadiya):

i just put that there to avoid confusion :)

OpenStudy (anonymous):

so then it would be.... a8=28+(8-1) x (-3) a8= 28=9*-3?

OpenStudy (diyadiya):

Yes Now tell me what is 9*-3 ?

OpenStudy (anonymous):

-27

OpenStudy (diyadiya):

Yes so 28+ (-27) =?

OpenStudy (anonymous):

55?

OpenStudy (anonymous):

1

OpenStudy (diyadiya):

Yes!

OpenStudy (anonymous):

a8=1 ? ? ?

OpenStudy (diyadiya):

Yes!

OpenStudy (anonymous):

a1=47, d=-5 , FIND a7

OpenStudy (diyadiya):

Try

OpenStudy (anonymous):

a7=47+(7-1) x (-5) ????

OpenStudy (diyadiya):

yup

OpenStudy (anonymous):

a7= 47+6*-5 47+-30=17 a7=17 ? ? ?

OpenStudy (diyadiya):

YESS you're doing great :D

OpenStudy (diyadiya):

How many more left?

OpenStudy (anonymous):

9 :\

OpenStudy (diyadiya):

ok you can do it ,i'll check if its right!

OpenStudy (anonymous):

ok a1=12 d=-7 FIND a12 a12=12+(12-1) x (-7) a12=12+11*-7 a12=12+-77 a12= -65 ?

OpenStudy (diyadiya):

RIGHT :) Next

OpenStudy (anonymous):

a1=18 d=-9 FIND a13 a13=18+(13-1) x (-9) a13=18+12*-9 a13= 18+-108 13= -90 ?

OpenStudy (diyadiya):

Right :D Next

OpenStudy (anonymous):

I NEED YOUR HELP ON THIS ONE !!!!!!!!!! 14,17,20,23,26,29..... FIND a13

OpenStudy (diyadiya):

ok the first number is a1

OpenStudy (anonymous):

is it 50????

OpenStudy (diyadiya):

how?

OpenStudy (anonymous):

you add 3 to get all of the shown numbers and i added 3 , 13 times and i had got 50

OpenStudy (diyadiya):

actually d is the common difference so to find the common difference we find out d (subtract first number from second number) so here a1 =14 d=17-14=3 so d=3 now how do you find a13?

OpenStudy (apoorvk):

Mammoth patience o.O Kudos.

OpenStudy (anonymous):

Not sure... but i REALLY neeed help on the next ones... t1=15, t2=21 t,20=?

OpenStudy (diyadiya):

Did you get the previous one ?

OpenStudy (anonymous):

No

OpenStudy (apoorvk):

Wanna continue here?

OpenStudy (diyadiya):

she posted new question

OpenStudy (diyadiya):

Ah shes here

OpenStudy (apoorvk):

There? @jordan199 ?